OpenStudy (anonymous):

find the absolute value of the complex number -8+3i

OpenStudy (anonymous):

The absolute value of a complex number is usually called the modulus. You compute it by finding the distance to the origin on the complex plane, basically the distance formula. In this case\[\left| -8+3i \right|=\sqrt{(-8)^2+3^2}=\sqrt{ 73}\]